OpenStudy (anonymous):

We toss 3 coins 100,000 times. How many times do we expect to see "Exactly 1 Head" appearing? A.)12,500 B.)37,500 C.)3,750 D.)50,000

OpenStudy (campbell_st):

toss 3 coins once and the probability of 1 head is 3/8 so from 1000000 tosses the expected number of 1 head is 3/8 x 1000000 = just evaluate for the answer

OpenStudy (anonymous):

outcomes = 2^3 = 8 i.e HHH, HHT ,HTT, HTH, THH, THT, TTH, TTT This means P(1head) = 1/8 from 100,000 tosses it becomes 1/8 x 100,000 = 12,500

OpenStudy (anonymous):

i did the problem the way campbell told me to and i got 100% , i dont think your answer is correct gita :/

OpenStudy (anonymous):

thanks both of you for taking the time to help me out though

OpenStudy (campbell_st):

hope it helps.

OpenStudy (anonymous):

you welcome